OpenStudy (anonymous):

Write an equation of a line whose y-intercept is 2 and whose slope is -1. A. y = -x + 2 B. y = x -1 C. y = -x -2 D. y = x + 1

OpenStudy (anonymous):

y=mx=b

OpenStudy (anonymous):

y=mx+b

OpenStudy (helder_edwin):

this problem is similar to one you already proposed. you should be able to do it on your own

OpenStudy (helder_edwin):

see what @HillDP just wrote

OpenStudy (anonymous):

b is the intercept m is the slope

OpenStudy (anonymous):

yea but i dont remember whhat gets plged into what after that i knw what to do

OpenStudy (anonymous):

a.y=-x+2

OpenStudy (anonymous):

*plugged

OpenStudy (shane_b):

Since @DHASHNI posted the answer, do you at least know why that's correct?

OpenStudy (anonymous):

no i still dont knw where to plug everything in

OpenStudy (anonymous):

you can do it

OpenStudy (shane_b):

You only have two values to plug in: the slope (which is m) and the intercept (which is b).

OpenStudy (anonymous):

b is the intercept m is the slope just substitute the values

OpenStudy (anonymous):

thanks thats what i needed to knw @Shane_B

OpenStudy (shane_b):

Good luck!

OpenStudy (anonymous):

@Shane_B : i do knw it y=mx+c is the general format of a line where m is the slope and c is the y-intercept

OpenStudy (shane_b):

I know you know...that question was meant for @suckatmath17. You posted the answer while we were trying to explain it to her.